# FastCuRL: Curriculum Reinforcement Learning with Stage-wise Context Scaling for Efficient Training R1-like Reasoning Models

### v5

@@ 5 Conclusion
We investigate how the model’s context length and the complexity of the training dataset influence the training process of R1-like reasoning models. Motivated by our findings, we propose FastCuRL, a simple yet effective curriculum reinforcement learning framework incorporating a stage-wise context scaling strategy. This framework is designed to accelerate the training efficiency and improve the model’s long CoT reasoning capabilities. Experimental results demonstrate that FastCuRL-1.5B-Preview achieves better performance and reduces computational resource consumption by more than 50%, with all training phases efficiently executed using a single node with 8 GPUs.
@@ 6 Limitations
Due to limited resources, this paper verifies the effectiveness of the proposed method, FastCuRL, only on a 1.5B language model. Generally, validating its effectiveness on models of varying sizes is a worthwhile direction for future research. Furthermore, in this paper, we investigate the influence of using complexity-aware training data by employing the simplest separation method to validate the efficacy of separating the training data by complexity, and achieves significant results. If more sophisticated separation methods were adopted, achieving even more promising results might be possible.
Training over multiple stages, rather than in a single training stage, involves more than changes in parameters like context length; it also fundamentally alters the reference policy. In a multi-stage training strategy, the KL penalty imposed by the reference policy on the model is gradually relaxed, which allows the trained model to explore a broader range of solutions. Delving into dynamic control of context lengths or implementing a dynamic KL penalty may be valuable directions.

### v6

@@ 5 Conclusion
We investigate how the model’s context length and the complexity of the training dataset influence the training process of R1-like reasoning models. Motivated by our findings, we propose FastCuRL, a simple yet effective curriculum reinforcement learning framework incorporating a stage-wise context scaling strategy. This method is designed to accelerate the training efficiency and improve the model’s long CoT reasoning capabilities. Experimental results demonstrate that FastCuRL-1.5B-Preview achieves better performance and reduces computational resource consumption by more than 50%, with all training phases efficiently executed using a single node with 8 GPUs.
@@ 6 Limitations
In this paper, we verify the proposed methodology exclusively on a 1.5B parameter language model. Extending the proposed methodology to models of diverse scales represents a critical avenue for future investigation. Moreover, while this work demonstrates the substantial impact of complexity-aware data selection through a relatively straightforward stratification approach, the observed performance gains suggest that more sophisticated data selection methodologies could potentially yield superior outcomes. Meanwhile, this observation motivates several research directions, including the development of adaptive context length and dynamic KL penalty controlling mechanisms, both of which could further optimize the balance between exploration and convergence in preference learning.